• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 09:33:34 on June16,2010

    Entry number 326900

    Followups:

    keyword=Change in respose of det 3 and 4

    During run 4594 response of detector 3 changed after event 51000. Beam current 
    is same and there is no halog for any change in high voltage of detectors. Is 
    it because of target?
    


    FIGURE 1